# Statistical Validation of Contagion Centrality in Financial Networks

### v1

@@ 2 Network Centrality
Though we constructed the above with a static network, we note that both the adjacency matrix and the weight vector for the node-level LCM can be time-varying. This will be used in the financial case studies of Section 4.
@@
System-level centrality is defined as the total sum of the node-level centralities $C$. This value provides a useful proxy for the total systemic risk captured in financial data.
@@ 3 Statistical Analysis
We go beyond traditional centrality calculations by deriving the asymptotic distribution of the LCM, enabling us to compute confidence intervals and perform statistical tests. As we calibrate the adjacency matrix as the OLS estimates, we recall this construction for a generic multivariate model. Given two matrices $X,Y\in\mathbb{R}^{T\times M}$, where $T$ represents the number of time steps and $M$ denotes the number of variables, and an iid noise term $\mathcal{E}\in\mathbb{R}^{T\times M}$, we can adopt a linear regression framework as below:
$\begin{gathered}Y=XA+\mathcal{E}\,,\\ \hat{A}=(X^{\top}X)^{-1}X^{\top}Y\,,\end{gathered}$ (4)
where $A$ is the true coefficient matrix that governs the relationships between the elements of $X$ and $Y$. On the other hand, $\hat{A}$ represents the estimated coefficient matrix, which is obtained from the regression analysis.
3.1 Distribution of the LCM
In practice, a discrepancy between the true adjacency matrix $A$ and its estimated counterpart $\hat{A}$ potentially exists. We denote this difference as $\Delta$, such that $\hat{A}=A+\Delta$. Taking advantage of the Taylor series expansion, we can express the LCM in terms of the discrepancy as follows:
@@ 3.1 Distribution of the LCM
Theorem 3 (LCM Distribution).
Assume the data follows a linear model. Let $\rho_{lk}$ denote the correlation coefficient between time series $l,k$. Let $\sigma_{k}$ denote the standard deviation of time series $k$. Define $Q:=\lim_{T\to\infty}T^{-1}X^{\top}X$ (with limit taken in probability as the number of data points grows to infinity). Set $\Sigma_{lk}:=\rho_{lk}\sigma_{l}\sigma_{k}Q^{-1}$ for each $l,k$. Then for every $i,j$ we have the following asymptotic distribution for the pair-level LCM:
@@
$\qquad\sim N\left(0,\left[(I-A)^{-1}_{\cdot j}\right]^{\top}\left[(I-A)^{-1}_{i\cdot}\Sigma_{lk}\left[(I-A)^{-1}_{i\cdot}\right]^{\top}\right]_{lk}(I-A)^{-1}_{\cdot j}\right)\,.$
This result follows from the normality of the error terms $\Delta_{\cdot j}\sim N(0,\Sigma_{jj})$ with covariance $\Sigma_{ij}$ between $\Delta_{\cdot i},\Delta_{\cdot j}$. The detailed proof of Theorem 3 is given in the Appendix A.1.
@@
Consider the same setting as in Theorem 3. The node-level LCM, defined in Equation 3, asymptotically follows the normal distribution:
@@ 3.2 Hypothesis Testing
In this section, we aim to explore a potential application of the node-level LCM distribution as outlined in Corollary 4. Specifically, we propose a statistical test to determine whether the node $i$ LCM significantly exceeds zero, i.e., the node contributes significantly to the propagation of a shock within the network.
@@
This result directly follows from Corollary 4 the Central Limit Theorem and Slutsky’s Theorem. The formal proof of Corollary 5 is given in Appendix A.3.
@@ 3.3 Numerical Validation
To evaluate the validity of the LCM, i.e., alignment between the theoretical and empirical distributions and their true counterparts, we conduct a case study using simulated data. We generate a dataset of 3 time series with iid noise terms and length of 600. The empirical analysis involves 10,000 simulations to obtain a robust representation. Subsequently, we calculate the mean and variance of the theoretical and empirical distributions of Leontief centrality measure and compare these distributions with the true values of the centrality.
Figure 2 presents the node-level LCM distributions. A close examination of the figures reveals a strong alignment between all three sets of values. The true centrality value falls within the 95% confidence interval calculated from the theoretical distribution. We note that the theoretical distribution may exhibit a slight bias due to the theoretical distribution’s mean being estimated from one realization of the adjacency matrix (see Figure 2(b)).
@@
Figure 2: Distribution of node-level LCM.
@@
Figure 3: QQ plot of node-level LCM comparing the theoretical and empirical distribution.
@@ 4 Financial Case Studies
In this section we present two case studies in order to demonstrate the financial implications of our framework. In the first case study, we utilize CDX spread data to construct networks, from which we analyze the LCM for a spanning history of over 11 years. In the second one, we leverage high-frequency equity tick data to capture the information in the stock market. We highlight the years with turmoils to demonstrate how the LCM reflects market behavior.
@@ 4.1 Credit Default Swap Index Market Contagion
In most of the existing literature, financial networks are typically constructed based on contracts between institutions, particularly in the banking sector, where a core of large banks are connected to peripheral ones. Numerous studies have explored this core-periphery network structure, mostly in the context of interbank lending (Battiston et al., , 2015; Bardoscia et al., , 2016; Bichuch and Feinstein, , 2018; Blasques et al., , 2018; Farboodi, , 2021). However, only a limited number of studies have delved into derivative markets, specifically Credit Default Swaps (CDS) (Brunnermeier et al., , 2013; Blasques et al., , 2016).
Our dataset is obtained from Bloomberg, breaking down Markit’s Investment Grade (IG) and High Yield (HY) CDX spread data [^1] For more information on these indices, we refer the interested reader to Appendix B. into sectors based on the Global Industry Classification Standard (GICS) [^2] The breakdown is at level 1 and excludes real estate.. This breakdown yields 10 sectors for each grade, totaling 20 time series. These indices have been reporting since January 1, 2011. Our analysis spans from this starting date to August 31, 2023. We use a one-year time window, sliding it monthly, resulting in 146 time windows. This approach allows us to assess the contagion risk across different periods, enabling a deeper understanding of the derivative market’s network’s dynamics over time.
To make the time series stationary and reduce variability, we compute the daily log return of the time series $r_{i,t}=log(\frac{s_{i,t+1}}{s_{i,t}})$ where $s_{i,t}$ is the CDX spread of the sector $i$ at time $t$. Using the framework with 1-day lag, we build our network for each time window. Afterwards, we calculate the LCM for the system. For simplicity, we perform the analysis and statistical tests based on the unweighted centrality setting ($\bm{w}=\bm{1}$).
@@
Figure 4: CDX network. Left February 2019-2020 Right March 2019-2020
@@
In Figure 5, we present the system-level Leontief centrality measure as a moving average across different time windows. The visualization indicates the overall level of contagion risk and sector instability, containing indices from both high yield and investment grade sectors. One key insight derived from the figure is the surge in LCM during and after the Covid-19 pandemic, indicating a high contagion risk during this period (Yu et al., , 2021). Comparing the time window from March 2019 to March 2020 with the preceding years, the measure obtains a value of about five times greater, reaching its peak in late 2020. Remarkably, when March is added to the time window, a sharp increase in LCM is triggered instantly, while it gradually starts to drop close to the end of 2020. These dynamics are primarily influenced by several events related to the outbreak of Covid-19 in early 2020 leading to a widespread economic disruptions. The uncertainty triggered significant market volatility with concerns about the economic fallout, indicating a higher perceived risk. Additionally, there were vulnerabilities specific to sectors that were captured in the CDX data. Afterwards, the US government and Federal Reserve responded with monetary policy interventions and fiscal stimulus in order to support financial markets. Signs of recovery were shown also when vaccines were rolled out, as optimism grew regarding the potential economic rebound. Economic indicators improved and reduced the uncertainty about the long-term pandemic’s impact. All these led to the gradual decline of risk and hence the LCM measure.
@@
We also notice a ramp up upon including mid-2022 data. This is mainly due to the increasing interest rates by Federal Reserve and some bankruptcy, merger and near-default events. It is worth mentioning that the elevated LCM in late 2012 may be due to the lingering effects of the Great Recession and the 2011 European sovereign debt crisis. This analysis highlights the measure’s robustness to economic events and interventions, showing its ability to capture the sector instability and contagion risk.
@@
We plot the leading eigenvalue and thresholded degree centrality alongside LCM in Figure 5. We see that the leading eigenvalue exhibits some level of noise, displaying marginal increase during Covid-19 period. Towards the end of the date range, this value maintains a consistent trend without any notable elevation. Similarly, the degree centrality line appears rather stable throughout the figure. This suggests that both leading eigenvalue and degree centrality struggle to capture the contagion risk effectively. For more in-depth analysis, Table 1 shows the percentage changes in these three metrics during the two periods of financial stress.
Figure 5: Moving average of CDX LCM, degree centrality and leading eigenvalue
@@
Table 1: Percentage change of centrality measures and leading eigenvalue during two periods of financial stress
Date LCM Degree Centrality Leading Eigenvalue
2020-2021 4.78283 0.07169 0.23087
2022-2023 5.52444 0.00096 -0.00312
@@
Figure 6 shows LCM over time for the Consumer Discretionary sector. We display the unvalidated LCM along with the 97.5% confidence interval. When the confidence interval intersects the zero line, the validated LCM is considered zero. In time window March 2019 to 2020, a notable increase in both investment grade and high yield CDX within the sector becomes evident. This time period coincides with the beginning and the outbreak of Covid-19 pandemic in U.S., disrupting economic activities across multiple sectors. Moreover, the lockdowns, social distancing and job losses affected the consumer behavior, leading to decreased spending on non-essential expenses such as leisure, travel and luxury items.
@@
On top of this, the global spread of Covid-19, which had already affected China, disrupted supply chains worldwide. Disruptions in the production and distribution of goods created challenges for companies within the discretionary sector, hindering their access to raw materials and logistics. Consequently, many businesses in non-essential retail sector either closed or experienced a major decline in sales. Hence, due to the pandemic, consumer behavior shifted, supply chain systems had disruptions and economic uncertainties escalated. Combination of these factors resulted into a notable increase in the contagion risk and vulnerability in the U.S. consumer discretionary sector. For a comprehensive view of sector-specific figures of CDX LCM, we refer the interested reader to Figure 10 in the Appendix.
@@
Figure 6: Consumer Discretionary sector CDX LCM Left Investment grade Right High yield
4.2 Equity Tick Data

### v2

@@ 2.2 Financial Intuition of KB
Though we constructed the above with a static network, we note that both the adjacency matrix and the weight vector for the node-level KB can be time-varying. This will be used in the financial case studies of Section 4. For the purposes of this work, recall that the adjacency matrix is provides weighted propagation of shocks through the network. Notably, this adjacency matrix can be derived using various methodologies, e.g., correlations, partial correlations, a vector autoregression model, or linear causality networks.
@@
When aiming to isolate idiosyncratic components from time series, one approach is to account for the influence of common factors. Techniques like PCA or Singular Value Decomposition (SVD) or spectral reduction (Ricciardi et al., , 2023) can identify dominant structures associated with these factors. However, to preserve the uniqueness of individual series, it is more appropriate to filter directly using observed factors relevant to the data, such as those derived from economic or financial theory.
@@ 3 Statistical Analysis
We extend the scope of traditional centrality calculations by developing the asymptotic distribution of the KB measure. This advancement not only allows for the computation of confidence intervals, but also facilitates statistical hypothesis testing. By deriving the distributional properties of the KB, we provide a framework to evaluate the significance of contagion risk estimates. This contributes to a deeper understanding of the measure’s reliability and makes it possible to compare results across different networks or over time with a higher degree of statistical rigor.
3.1 Distribution of KB
In practice, a discrepancy between the true adjacency matrix $A$ and its estimated counterpart $\hat{A}$ potentially exists. We denote this difference as $\Delta$, such that $\hat{A}=A+\Delta$. For $T$ data points, taking advantage of the Taylor series expansion, we can express the KB in terms of the discrepancy as follows:
@@ 3.1 Distribution of KB
By applying the Delta method (Greene, , 2003), we are able to deduce the asymptotic distribution for the pair-level KB. Throughout this section we consider the shorthand notation where $\cdot j$ denotes the $j$th column and $i\cdot$ denotes the $i$th row, respectively.
Theorem 3 (KB Distribution).
Assume $\Delta_{\cdot j}\sim N(0,\Sigma_{jj})$ where $Cov(\Delta_{\cdot i},\Delta_{\cdot j})=\Sigma_{ij}$. Then for every $i,j$ we have the following asymptotic distribution for the pair-level KB:
@@
$\qquad\sim N\left(0,\left[(I-A)^{-1}_{\cdot j}\right]^{\top}\left[(I-A)^{-1}_{i\cdot}\Sigma_{lk}\left[(I-A)^{-1}_{i\cdot}\right]^{\top}\right]_{lk}(I-A)^{-1}_{\cdot j}\right)\,,$
The detailed proof of Theorem 3 is given in the Appendix A.1. This consideration of the pair-level KB distribution permits us to consider, also, the distribution of node-level KB.
@@
Consider the same setting as in Theorem 3. The node-level KB, defined in Equation node-level KB, asymptotically follows the normal distribution:
@@ 3.2 Hypothesis Testing
In this section, we aim to explore a potential application of the node-level KB distribution as outlined in Corollary 4. Specifically, we propose two statistical tests on the KB centrality to determine its reliability in practice. First, we introduce a statistical test to determine whether the node $i$ KB significantly exceeds zero, i.e., the node contributes significantly to the propagation of a shock within the network. A second hypothesis test examines whether the centrality values of two nodes differ significantly from each other.
3.2.1 Different from Zero
In this hypothesis test, we aim to determine whether the node-level centrality significantly differs from zero. This approach adds an additional layer of validation to the centrality measure, ensuring that the results are not merely driven by noise but are statistically meaningful.
@@ 3.2.1 Different from Zero
This result directly follows from Corollary 4, the Central Limit Theorem and Slutsky’s Theorem. The formal proof is given in Appendix A.3.
3.2.2 Different from Each Other
To compare different shock scenarios, specifically evaluating the contagion risk when node $i$ is shocked versus when node $j$ is shocked, the following hypothesis test can be carried out:
$\begin{gathered}H_{0}:(I-A)^{-1}_{i\cdot}\bm{w}-(I-A)^{-1}_{j\cdot}\bm{w}=0\,,\\ H_{1}:(I-A)^{-1}_{i\cdot}\bm{w}-(I-A)^{-1}_{j\cdot}\bm{w}\neq 0\,.\end{gathered}$ (3)
Corollary 6 (Test Statistic and Distribution).
Consider the setting of Theorem 3 with the test statistic:
$Z:=\frac{\sqrt{T}\left[(I-\hat{A})^{-1}_{i\cdot}\bm{w}-(I-\hat{A})^{-1}_{j\cdot}\bm{w}\right]}{\left(Var\left[(I-\hat{A})^{-1}_{i\cdot}\bm{w}\right]+Var\left[(I-\hat{A})^{-1}_{j\cdot}\bm{w}\right]-Cov\left[(I-\hat{A})^{-1}_{i\cdot}\bm{w},(I-\hat{A})^{-1}_{j\cdot}\bm{w}\right]\right)^{\frac{1}{2}}}\,$
where the variances can be obtained from Corollary 4 and the covariance structure is given by:
$Cov\left[(I-\hat{A})^{-1}_{i\cdot}\bm{w},(I-\hat{A})^{-1}_{j\cdot}\bm{w}\right]=\left[(I-\hat{A})^{-1}\bm{w}\right]^{\top}\left[(I-\hat{A})^{-1}_{i\cdot}\Sigma_{lk}\left[(I-\hat{A})^{-1}_{j\cdot}\right]^{\top}\right]_{lk}(I-\hat{A})^{-1}\bm{w}\,.$
Under the null hypothesis, $Z\sim N(0,1)$ asymptotically.
The proof follows identically to that of Corollary 5 and is given in Appendix A.3.
@@ 3.3 Numerical Validation
To evaluate the validity of the KB, i.e., alignment between the theoretical and empirical distributions and their true counterparts, we conduct a case study using simulated data. We generate a dataset of 3 linearly dependent time series with iid noise terms and length of 600. The empirical analysis involves 10,000 simulations to obtain a robust representation. Subsequently, we calculate the mean and variance of the theoretical and empirical distributions of the KB centrality measure and compare these distributions with the true values of the centrality.
Figure 2 presents the node-level KB distributions. A close examination of the figures reveals a strong alignment between all three sets of values. The true centrality value falls within the 95% confidence interval calculated from the theoretical distribution. We note that the theoretical distribution may exhibit a slight bias due to the theoretical distribution’s mean being estimated from one realization of the adjacency matrix (see Figure 2(b)).
@@
Figure 2: Distribution of node-level KB.
@@
Figure 3: QQ plot of node-level KB comparing the theoretical and empirical distribution.
@@ 4 Financial Case Studies
In this section we present two case studies in order to demonstrate the financial implications of our framework. In the first case study, we utilize CDX spread data to construct networks, from which we analyze the KB for a spanning history of over 11 years. In the second one, we leverage high-frequency equity tick data to capture the information in the stock market. We highlight the years with turmoils to demonstrate how the KB reflects market behavior. For simplicity, we obtain the adjacency matrix from a Vector Autoregressive Linear System considering one lag (VAR(1)). For more details refer to Appendix B.
Remark 5.
Herein, to overcome parameter instability of the adjacency matrix when calibrated from a vector autoregressive model, we apply moving average smoothing to all following results. We refer the interested reader to, e.g., Diebold and Yilmaz, (2009); Kilian and Lütkepohl, (2018); Sudarshan Kumar and Chakrabarti, (2022); Diebold and Yilmaz, (2023) on difficulties with the use of vector autoregressive models.
@@ 4.1 Credit Default Swap Index Market Contagion
In most of the existing literature, financial networks are typically constructed based on contracts between institutions, particularly in the banking sector, where a core of large banks are connected to peripheral ones. Numerous studies have explored this core-periphery network structure, mostly in the context of interbank lending (Upper, , 2011; Battiston et al., , 2015; Bardoscia et al., , 2016; Bichuch and Feinstein, , 2018; Blasques et al., , 2018; Farboodi, , 2021). However, only a limited number of studies have delved into derivative markets, specifically Credit Default Swaps (CDS) (Brunnermeier et al., , 2013; Blasques et al., , 2016).
Our dataset is obtained from Bloomberg, breaking down Markit’s Investment Grade (IG) and High Yield (HY) CDX spread data [^3] For more information on these indices, we refer the interested reader to Appendix C. into sectors based on the Global Industry Classification Standard (GICS) [^4] The breakdown is at level 1 and excludes real estate.. This breakdown yields 10 sectors for each grade, totaling 20 time series. These indices have been reporting since January 1, 2011. Our analysis spans from this starting date to August 31, 2023. We use a one-year time window, sliding it monthly, resulting in 146 time windows. This approach allows us to assess the contagion risk across different periods, enabling a deeper understanding of the derivative market’s network’s dynamics over time.
To make the time series stationary and reduce variability, we compute the daily log return of the time series $r_{i,t}=log(\frac{s_{i,t+1}}{s_{i,t}})$ where $s_{i,t}$ is the CDX spread of the sector $i$ at time $t$. Using the framework with 1-day lag, we build our VAR(1) network from these daily returns for each time window. Afterwards, we calculate the KB for the system. For simplicity, we perform the analysis and statistical tests of Section 3.2.1 based on the unweighted centrality setting ($\bm{w}=\bm{1}$).
@@
Figure 4: CDX network. Left February 2019-2020 Right March 2019-2020. The blue colored nodes are Energy and Consumer Discretionary, IG and HY
@@
In Figure 5, we present the system-level KB centrality measure as a moving average across different time windows. The visualization indicates the overall level of contagion risk and sector instability, containing indices from both high yield and investment grade sectors. One key insight derived from the figure is the surge in KB during and after the Covid-19 pandemic, indicating a high contagion risk during this period (Yu et al., , 2021). Comparing the time window from March 2019 to March 2020 with the preceding years, the measure obtains a value of about five times greater, reaching its peak in late 2020. Remarkably, when March is added to the time window, a sharp increase in KB is triggered instantly, while it gradually starts to drop close to the end of 2020. These dynamics are primarily influenced by several events related to the outbreak of Covid-19 in early 2020 leading to a widespread economic disruptions. The uncertainty triggered significant market volatility with concerns about the economic fallout, indicating a higher perceived risk. Additionally, there were vulnerabilities specific to sectors that were captured in the CDX data. Afterwards, the US government and Federal Reserve responded with monetary policy interventions and fiscal stimulus in order to support financial markets. Signs of recovery were shown also when vaccines were rolled out, as optimism grew regarding the potential economic rebound. Economic indicators improved and reduced the uncertainty about the long-term pandemic’s impact. All these led to the gradual decline of risk and hence the KB measure.
@@
We also notice a ramp up upon including mid-2022 data. This is mainly due to the increasing interest rates by Federal Reserve and some bankruptcy, merger and near-default events. It is worth mentioning that the elevated KB in late 2012 may be due to the lingering effects of the Great Recession and the 2011 European sovereign debt crisis. This analysis highlights the measure’s robustness to economic events and interventions, showing its ability to capture the sector instability and contagion risk.
@@
We plot the leading eigenvalue and thresholded degree centrality alongside KB in Figure 5 for the same adjacency matrix. We see that the leading eigenvalue exhibits some level of noise, displaying marginal increase during Covid-19 period. Towards the end of the date range, this value maintains a consistent trend without any notable elevation. Similarly, the degree centrality line appears rather stable throughout the figure. Though not displayed, we also calculated the betweenness and closeness centralities; these exhibited similar patterns to degree centrality. This suggests that leading eigenvalue and the other three centralities struggle to capture the contagion risk effectively.
Figure 5: Moving average of CDX KB, degree centrality and leading eigenvalue
In Figure 6, we compare the KB to its unvalidated counterpart and the DebtRank measure (Battiston et al., 2012b, ) taken with respect to the same adjacency matrix. Unvalidated counterpart leaves out the hypothesis test in Equation 2. DebtRank (DR) is a feedback centrality measure that distresses one node at a time and quantifies the propagation of systemic risk throughout the network. At each time step, nodes are classified as undistressed, distressed, or inactive.
From mid-2013 through the end of 2019, both the unvalidated KB and DR show a gradual rise, with DR beginning its increase earlier in 2013 and the unvalidated KB starting to climb in 2015. However, this period, from 2013 to 2019, was marked by relative financial stability, [^5] https://www.forbes.com/advisor/investing/bull-market-history/#:~:text=Bull%20Market%20of%202009%2D2020%3A%20The%20Great%20Recession%20Recovery with no significant systemic shocks, making the steady rise in both measures noticeable. However, this increase may reflect false alarms due to the lack of statistical validation. In contrast, the validated KB behaves differently. While some noise is present, there is no significant upward trend during this stable period. This highlights the importance of validation in filtering out spurious signals which allows the KB to reduce the likelihood of false positives.
During the Covid period, all three measures rise, capturing the financial turbulence. However, when examining DR, the increase in early 2021 is nearly identical in magnitude to the peak observed in September 2019, both reaching 0.8. The 2019 peak occurred without any clear signs of systemic instability, emphasizing the challenge of interpreting unvalidated metrics.
All three measures experience a decline after the end of 2021, followed by an upward trend as the financial system reacts to the risks introduced by rising interest rates. For more in-depth analysis, Table 1 shows the percentage changes in these measures during the two periods of financial stress.
Figure 6: Moving average of CDX KB, not validated KB and DebtRank
@@
Table 1: Relative change of centrality measures and leading eigenvalue during two periods of financial stress
Date KB Degree Centrality Leading Eigenvalue DebtRank Unvalidated KB
March 2019 - March 2020 5.327 0.123 0.025 -0.387 0.763
March 2022 - March 2023 7.260 0.003 0.077 3.773 8.533
@@
Figure 7 shows KB over time for the Consumer Discretionary sector. We display the unvalidated KB along with the 97.5% confidence interval. When the confidence interval intersects the zero line, the validated KB is considered zero. In time window March 2019 to 2020, a notable increase in both investment grade and high yield CDX within the sector becomes evident. This time period coincides with the beginning and the outbreak of Covid-19 pandemic in U.S., disrupting economic activities across multiple sectors. Moreover, the lockdowns, social distancing and job losses affected the consumer behavior, leading to decreased spending on non-essential expenses such as leisure, travel and luxury items.
@@
On top of this, the global spread of Covid-19, which had already affected China, disrupted supply chains worldwide. Disruptions in the production and distribution of goods created challenges for companies within the discretionary sector, hindering their access to raw materials and logistics. Consequently, many businesses in non-essential retail sector either closed or experienced a major decline in sales. Hence, due to the pandemic, consumer behavior shifted, supply chain systems had disruptions and economic uncertainties escalated. Combination of these factors resulted into a notable increase in the contagion risk and vulnerability in the U.S. consumer discretionary sector. For a comprehensive view of sector-specific figures of CDX KB, we refer the interested reader to Figure 11 in the Appendix.
@@
Figure 7: Consumer Discretionary sector CDX KB Left Investment grade Right High yield
4.2 Equity Tick Market Contagion